Interactive Solution 18.43 provides a model for problems of this kind. A small object has a mass of 3.0 x 10 there is an electric field, when released, the object experiences an acceleration of 2.6 10, m/s2 in the di relative to the +x axis. charge of 4 rection of the +x axis. Determine the elect kg and a S4C. It is placed at a certain spot where including sign, Number
Answer #1

as per equilibrium of forces

ma = qE

3*10^-3* 2.6*10^3 = 45*10^-6* E

E = 1.73*10^5 N/C

=======

Comment in case any doubt, will reply for sure.. Goodluck